# Optimizing Integrated Terrestrial and Non-Terrestrial Networks Performance with Traffic-Aware Resource Management

### v1

@@ III Problem Formulation
Our objective is to develop a framework, which simultaneously balances the load amongst the MBS and reduces the total power consumption in the terrestrial network by adjusting the resource distribution in response to fluctuations in network traffic. Specifically, our goal is to achieve proportional fair resource allocation by optimizing the network SLT [11]: in fact, the nature of the logarithmic cost function discourages each MBS to allocate disproportionate resources to a single UE. Let us denote as $\varepsilon$ the fraction of the bandwidth allocated to the LEO satellites; then, the bandwidth allocated for an MBS $j$ can be written as:
@@ IV Designed Solution
In this section, we first present BLASTER, the framework proposed to address the optimization problem outlined in equations (III)-(III). Then, we design a low-complexity heuristic based on the special characteristics of the problem, and provide a comparison for BLASTER w.r.t. performance state-of-the-art benchmarks. Finally, we provide an analysis of the complexity of both solutions.
@@ IV-A BLASTER
       Compute: $p{(s+1)}=\Bigr{[}\hat{p}(s)\Bigr{]}_{\tau}^{p^{\rm MAX}}$               (26);
@@ IV-A1 Utility optimization under fixed transmit power
Note that, since in (17) both the MBS transmit power and the noise power scale linearly with the bandwidth, $\gamma_{ij}$ is unaffected by the bandwidth split between satellite and terrestrial network.
@@ IV-A2 Transmit power optimization under fixed association
A closed-form solution to this problem, known as block-soft thresholding [28, Sec. 6.5.1] is expressed as:
@@ IV-B Heuristic development
  • UE Association: To maximize the number of terrestrial MBSs that can be switched off during low-traffic hours, it is desirable to maximize the share of the traffic served by the non-terrestrial tier. Hence, the proposed heuristic associates each UE that has an RSRP larger than $RSRP_{\min}$ to the satellite. For each remaining UEs that is yet to be associated to a MBS, we collect the perceived RSRPs from all the MBSs that are providing an RSRP greater than $\geq RSRP_{\rm min}$. Then, we average these RSRP measurements and rank each UE based on the calculated average from worst to best. Then, starting from the UE with the poorest average, we associate it with the MBS providing the largest throughput with the aim of maximizing the SLT.
    In the case of a high-traffic scenario, the satellite merely acts as an umbrella over the ground network to provide service for UEs that have no service, and potentially facilitate the distribution of the load. In this regard, the UE association process is quite similar to the low-traffic scenario as we rank each UE based on their average perceived RSRP and associate to the MBS providing the best throughput [^2] Note that the association is done only once for each UE and does not change even if the associated MBS is not providing the best throughput available anymore..
@@
  • MBS shutdown and power control: In low-traffic scenario, the proposed heuristic shuts down all MBSs serving less UEs than a set threshold $T_{UE}$, provided that these UEs can be successfully handed over to neighbouring MBSs. MBSs. We also reduce the MBS transmit power as long as the coverage constraint (III) is satisfied for the connected UEs (feasible region computed in (25)). On the other side, for high traffic, we only shut down inactive MBSs and reduce the transmit power similarly to the low-traffic scenario.
@@
      // For remaining active BSs:
@@ V Simulation Results & Analysis
In this section, we assess the performance of the proposed solution under different traffic conditions. Specifically, in Sec. V-A, we describe the simulation settings and the baseline algorithms, and in Sec. V-B we discuss the impact of the regularization parameter $\lambda$, introduced in (III), on the targeted performance metrics. In Sec. V-D, we describe how the algorithms under investigation leverage the satellite network resources to offload the terrestrial network. Finally, in Secs. V-E and V-F, we analyse the performance of the solutions under tests in terms of network sum-log throughput and terrestrial network energy consumption, respectively.
@@ V-A Simulation settings and benchmarks
The area of study is approximately $2500$ $\mathrm{km}^{2}$, which matches the beam diameter of a LEO satellite [29], and includes both an urban and a rural area, as depicted in Figure 1. In this area, we model the traffic variations by changing the number of active UEs in the network. Precisely, the number of UEs deployed in the network scales according to the daily downlink traffic load pattern presented in [30]. The number of UEs decreases during the night, until it reaches its minimum of $400$ at $5$ AM, and increases in high traffic, reaching up to $10000$ at $8$ PM. The UEs are deployed uniformly across the entire study area, with a higher density of deployment in the urban area relatively to the rural one. The terrestrial MBSs in both urban and rural areas are deployed in a hexagonal grid layout [31], with a higher density of MBSs in the urban area. In our simulation setting, we consider that $80\%$ of the UEs are indoor [23].
@@ V-A1 Satellite deployment
In our study, we consider a LEO satellite constellation employing earth-fixed beams [24, Section 4.6], implying that the satellites use beam pointing mechanisms (mechanical or electronic steering feature) to compensate for their mobility and cover a given fixed region. Specifically, we assume in our study that a satellite has a seven beam configuration, as depicted in red in Figure 2, and that at any given time, a satellite in the constellation is able to serve the region of interest.
As stated in Section II, the elevation angle of a satellite is an important parameter, which impacts the quality of the signal perceived from the satellite. Taking this into account, we embrace the mobility of the LEO satellites by assessing network performance at three distinct positions (P1, P2, P3), as also illustrated in Figure 2:
@@ 1. P1)
    We make the first performance assessment when the satellite is located $50$ kms from the center of the area of study, with an elevation angle (for a UE located at the center of the central beam) approximately equal to $84^{\circ}$.
@@ 3. P3)
    Similarly to the first measurement, we make the assessment when the satellite is $50$ kms away from the center of the area of study.
@@ V-A2 Benchmarks
To compare and assess the performance of both developed algorithms, we introduce the two following benchmarks: The 3GPP-TN configuration includes only a terrestrial network operating on a bandwidth of $10$ MHz. In contrast, the 3GPP-NTN configuration includes a satellite network overlaying the terrestrial one. In this case, the bandwidth is divided according to the 3GPP suggestions in [29], with $30$ MHz assigned to the satellite tier and $10$ MHz to the terrestrial one. In both configurations, the UEs associate to the MBSs based on the max-RSRP rule, with no DL transmit power optimization or MBS shutdown implemented. The most important simulation parameters are listed in Table I following [32, 24, 29, 23, 33, 34].
@@
Terrestrial Max Tx Power per RE $p_{j}^{MAX}$ [33] 17.7 $17.7\text{ }\mathrm{dBm}$
Satellite Max Tx Power per RE $p_{j}^{MAX}$ [29] 15.8 $15.8\text{ }\mathrm{dBm}$
@@
TABLE I: Simulation parameters.
@@ V-B Optimization of the regularization parameter
As specified in Section III, $\lambda$ is a regularization parameter, which allows to control the trade-off between maximizing network SLT and minimizing the terrestrial network power consumption (see Eq. (III)). Indeed, a high value of $\lambda$ leads BLASTER to increase the number of shutdown MBSs, as the priority becomes to reduce the terrestrial network power consumption. Conversely, a low value of $\lambda$ indicates that the focus is on improving the SLT, by balancing the load and ensuring a proportional fair resource allocation. To achieve this goal, in our study, we set $\lambda$ inversely proportional to the number of UEs, $K$, in the network as follows:
@@
where $\lambda_{\mathrm{max}}$ is the value of $\lambda$ when the UEs number is equal to its lowest, i.e., $K_{\mathrm{min}}$. This approach allows us to simplify the setting of the hyper-parameter $\lambda$, by fixing $\lambda_{\rm max}$ and using (28). Therefore, in the following, we investigate the impact of $\lambda_{\rm max}$ on the behaviour of BLASTER. In particular, we let $\lambda_{\rm max}$ take values in $\left\{10,1e3,1e7,1e10\right\}$, and study the corresponding distribution of the SLT during high and low traffic as well as that of the total terrestrial network power consumption.
In Figure 4, we see a clear trend, with larger values of $\lambda_{\rm max}$ leading to a downward shift of the SLT distribution during high traffic. In fact, we observe a gradual increase of the median value of the SLT in high traffic, as $\lambda_{\rm max}$ decreases. This is in line with our assumption that a small value of $\lambda$ moves the focus from the terrestrial network power consumption to improving the SLT. Specifically, in Figure 4, we see a $31\,\%$ increase of SLT by varying $\lambda_{\rm max}$ from the largest value ($\lambda_{\rm max}=1e10$) to the smallest one ($\lambda_{\rm max}=10$). Note that, during high traffic, the performance in terms of SLT stagnates, once $\lambda_{\rm max}$ starts crossing extremely high values. Conversely, in Figure 4, we notice that there is a limited change of the SLT, if the value of $\lambda_{\rm max}$ varies. In fact, in low traffic scenarios, our framework aims mainly to reduce the terrestrial network power consumption, i.e., maximising the SLT has limited importance.
To verify this analysis, Figure 5 plots the terrestrial network energy consumption for the distinct values of $\lambda_{\rm max}$, specified previously. We remark that the terrestrial network energy consumption increases throughout the day when the value of $\lambda_{\rm max}$ decreases. Indeed, a small value of $\lambda_{\rm max}$ hampers the transmit power optimization in (23), as each update becomes negligible, making it hard to effectively drive down the energy consumption of the terrestrial network.
Through these results, we are able to grasp the trade-off between reducing power consumption and balancing the network load efficiently. Indeed, we observe that limiting the value of $\lambda_{\rm max}$, i.e., $\lambda_{\rm max}=10$ and $\lambda_{\rm max}=1e3$, we only achieve, during high traffic hours, a minor gain in terms of network SLT of approximately $2\%$ compared to the other cases ($\lambda_{\rm max}=1e7$ and $\lambda_{\rm max}=1e10$), while the total power consumption surges by approximately $84\%$. Considering all of this, we fix $\lambda_{\rm max}=1e7$ as this value ensures a balanced performance in both SLT and terrestrial network energy consumption.
@@
V-C Complexity Analysis
In Figure 6, we display the operational complexity of both proposed frameworks BLASTER and HEURISTIC as derived in Section IV-C. We notice that the operational complexity of BLASTER is higher than HEURISTIC throughout the day. This is expected, as the latter framework is implemented using less complex operations compared to the former, which resorts to multiple gradient descent methods. In fact, we see an average decrease in operational complexity of approximately $21\%$ in the day, which underlines the simplicity of HEURISTIC compared to BLASTER. Moreover, we notice that the complexity follows a pattern similar to the traffic load for both frameworks, which corroborates with the formulas derived in IV-C. Indeed, a lower number of total UEs $K$ naturally leads to a lower complexity, as seen in Figure 6. On the opposite, a surge in traffic results in an increased complexity due to a higher number of operations needed to complete both frameworks.
@@ V-D Impact of the Satellite Network on Traffic Distribution
In this section, we look at the shifting role that the satellites play on the mobile network throughout the day. Figure 7 presents the daily profile of the fraction of UEs associated with the satellite network in the different frameworks under investigation. Also, we plot the hourly number of UEs deployed in the network, represented by the black dotted line. As we have previously underlined, the priority for the proposed BLASTER and HEURISTIC frameworks during low traffic is to reduce the terrestrial network power consumption. Hence, the satellite becomes a compelling option for offloading the terrestrial network and shutting down lightly loaded MBSs. Figure 7 shows that BLASTER and HEURISTIC achieve a $70\,\%$ and $500\,\%$ increase of the fraction of UEs associated to the satellite network in low-traffic hours as compared to the benchmark 3GPP-NTN. Also, we notice that the proportion of UEs associated to the satellite is far greater for HEURISTIC compared to BLASTER. This is due to the fact that HEURISTIC ensures that every UE that has a signal strength greater than $RSRP_{\rm min}$ is associated to the satellite, opposed to the more sophisticated BLASTER, who would still consider the available throughput before associating to the satellite. Conversely, during high traffic hours, we notice that the satellite network takes a less prominent role for both BLASTER and HEURISTIC, essentially acting as an umbrella, providing to UEs that would otherwise be out of coverage. Accordingly, most of of the total bandwidth is allocated to the terrestrial tier, which can efficiently serve more UEs than the non-terrestrial tier and can significantly improve the spatial reuse.

### v2

@@ III Problem Formulation
Our objective is to develop a framework, which simultaneously increases UE performance and reduces network energy consumption by adjusting the resource distribution between satellite and terrestrial MBSs in response to the hourly fluctuations in network traffic. Specifically, our goal is to achieve proportional fair resource allocation by optimizing the network SLT, while limiting the TN energy consumption: in fact, the nature of the logarithmic cost function discourages each MBS to allocate disproportionate resources to a single UE. Let us denote as $\varepsilon$ the fraction of the bandwidth allocated to the LEO satellites at a given hour of the day; then, the bandwidth allocated for an MBS $j$ can be written as:
@@ IV Designed Solution
In this section, we first present BLASTER, the framework proposed to address the optimization problem outlined in (III)-(III). Then, we design a low-complexity heuristic based on the special characteristics of the problem, and provide a comparison for BLASTER w.r.t. performance state-of-the-art benchmarks. Finally, we provide an analysis of the complexity of both solutions.
@@ IV-A BLASTER
       Compute: $p{(s+1)}=\Bigr{[}\hat{p}(s)\Bigr{]}_{\tau}^{p_{\rm max}}$                   (26);
@@ IV-A1 Utility optimization under fixed transmit power
Note that, since in (17) both the MBS transmit power and the noise power scale linearly with the bandwidth, $\gamma_{ij}$ is unaffected by the bandwidth split between satellite and TN.
@@ IV-A2 Transmit power optimization under fixed association
A closed-form solution to this problem, known as block-soft thresholding [38, Sec. 6.5.1] is expressed as:
@@ IV-B Heuristic development
  • UE Association: To maximize the number of terrestrial MBSs that can be switched off during low-traffic hours (00 AM - $7$ AM), it is desirable to maximize the share of the traffic served by the non-terrestrial tier. Hence, the proposed heuristic associates each UE that has an RSRP larger than $RSRP_{\min}$ to the satellite. For each remaining UEs that is yet to be associated to a MBS, we collect the perceived RSRPs from all the MBSs that are providing an RSRP greater than $\geq RSRP_{\rm min}$. Then, we average these RSRP measurements and rank each UE based on the calculated average from worst to best. Then, starting from the UE with the poorest average, we associate it with the MBS providing the largest throughput with the aim of maximizing the SLT.
    In the case of a high-traffic scenario, the satellite merely acts as an umbrella over the ground network to provide service for UEs that have no service, and potentially facilitate the distribution of the load. In this regard, the UE association process is quite similar to the low-traffic scenario as we rank each UE based on their average perceived RSRP and associate to the MBS providing the best throughput [^2] Note that the association is done only once for each UE and does not change even if the associated MBS is not providing the best throughput anymore due to the increased number of served UEs..
@@
  • MBS shutdown and power control: In low-traffic scenario, the proposed heuristic shuts down all MBSs serving less UEs than a set threshold $T_{UE}$, provided that these UEs can be successfully handed over to neighbouring MBSs. We also reduce the MBS transmit power as long as the coverage constraint (III) is satisfied for the connected UEs (feasible region computed in (25)). On the other side, for high traffic, we only shut down inactive MBSs and reduce the transmit power similarly to the low-traffic scenario.
@@
      // For remaining active base stations:
@@ V Simulation Results & Analysis
In this section, we assess the performance of the proposed solution under different traffic conditions. Specifically, in Sec. V-A, we describe the simulation settings and the baseline algorithms, and in Sec. V-B we discuss the impact of the regularization parameter $\lambda$, introduced in (III), on the targeted performance metrics. In Sec. V-D, we describe how the algorithms under investigation leverage the satellite network resources to offload the TN. Finally, in Secs. V-E and V-F, we analyse the performance of the solutions under tests in terms of network sum-log throughput and TN energy consumption, respectively.
@@ V-A Simulation settings and benchmarks
All the results and analyses presented in the following section have been conducted over a $24$-hour observation window, with snapshots of the network taken at the start of each hour, and generated using a custom-built system-level simulator, which adheres to 3GPP recommendations [39, 32, 40, 31, 41, 42]. The area of study is approximately $2500$ $\mathrm{km}^{2}$, which matches the beam diameter of a LEO satellite [40], and includes both an urban and a rural area, as depicted in Fig. 1. In this area, we model the traffic variations by changing the number of active UEs in the network. Precisely, the number of UEs deployed in the network scales according to the daily downlink traffic load pattern presented in [43]. The number of UEs decreases during the night, until it reaches its minimum of $400$ at $5$ AM, and increases in high traffic, reaching up to $10000$ at $8$ PM. The UEs are deployed uniformly across the entire study area, with a higher density of deployment in the urban area relatively to the rural one. The terrestrial MBSs in both urban and rural areas are deployed in a hexagonal grid layout [44], with a higher density of MBSs in the urban area. In our simulation setting, we consider that $80\%$ of the UEs are indoor [31].
@@ V-A1 Satellite deployment
In our study, we consider a LEO satellite constellation employing earth-fixed beams [32, Section 4.6], implying that the satellites use beam pointing mechanisms (mechanical or electronic steering feature) to compensate for their mobility and cover a given fixed region. Specifically, we assume in our study that a satellite has a seven beam configuration, as depicted in Fig. 2, and that at any given time, a satellite in the constellation is able to serve the region of interest in red.
As stated in Section II, the elevation angle of a satellite is an important parameter, which impacts the quality of the signal perceived from the satellite. Taking this into account, we embrace the mobility of the LEO satellites by assessing network performance at three distinct positions (P1, P2, P3), as also illustrated in Fig. 2:
@@ 1. P1)
    We make the first performance assessment when the satellite is located at the nadir point of the beam adjacent to our area of study, which is located at a horizontal distance of $50$ kms (as well as a total distance of approximately $603$ kms) from the center of the area of study, with an elevation angle (for a UE located at the center of the central beam) approximately equal to $84^{\circ}$.
@@ 3. P3)
    Similarly to the first measurement, we make the assessment when the satellite is $50$ kms away from the center of the area of study, at the nadir point of the next contiguous beam.
@@ V-A2 Benchmarks
To compare and assess the performance of both developed algorithms, we introduce the two following benchmarks: The 3GPP-TN configuration includes only a TN operating on a bandwidth of $10$ MHz. In contrast, the 3GPP-NTN configuration includes a satellite network overlaying the terrestrial one. In this case, the bandwidth is divided according to the 3GPP suggestions in [40], with $30$ MHz assigned to the satellite tier and $10$ MHz to the terrestrial one. In both configurations, the UEs associate to the MBSs based on the max-RSRP rule, with no DL transmit power optimization or MBS shutdown implemented. The most important simulation parameters are listed in Table II following [39, 32, 40, 31, 41, 42].
@@
Terrestrial Max Tx Power per RE ${p}_{\mathrm{max}}$ [41] 17.7 $17.7\text{ }\mathrm{dBm}$
Satellite Max Tx Power per RE ${p}_{\mathrm{max}}$ [40] 15.8 $15.8\text{ }\mathrm{dBm}$
@@
Satellite baseline energy consumption $E_{c}$ 500 $500\text{ }\mathrm{J}$
TABLE II: Simulation parameters.
@@ V-B Optimization of the regularization parameter
As specified in Section III, $\lambda$ is a regularization parameter, which allows to control the trade-off between maximizing network SLT and minimizing the TN energy consumption (see (III)). Indeed, a high value of $\lambda$ leads BLASTER to increase the number of shutdown MBSs, as the priority becomes to reduce the TN energy consumption. Conversely, a low value of $\lambda$ indicates that the focus is on improving the SLT, by balancing the load and ensuring a proportional fair resource allocation. To achieve this goal, in our study, we set $\lambda$ inversely proportional to the number of UEs, $K$, in the network as follows:
@@
where $\lambda_{\mathrm{max}}$ is the value of $\lambda$ when the UEs number is equal to its lowest, i.e., $K_{\mathrm{min}}$. This approach allows us to simplify the setting of the hyper-parameter $\lambda$, by fixing $\lambda_{\rm max}$ and using (28). Therefore, in the following, we investigate the impact of $\lambda_{\rm max}$ on the behaviour of BLASTER. In particular, we let $\lambda_{\rm max}$ take values in $\left\{10,1e3,1e7,1e10\right\}$, and study the corresponding distribution of the SLT during high and low traffic as well as that of the total TN energy consumption.
In Fig. 4, we see a clear trend, with larger values of $\lambda_{\rm max}$ leading to a downward shift of the SLT distribution during high traffic. In fact, we observe a gradual increase of the median value of the SLT in high traffic, as $\lambda_{\rm max}$ decreases. This is in line with our assumption that a small value of $\lambda$ moves the focus from the TN energy consumption to improving the SLT. Specifically, in Fig. 4, we see a $2\,\%$ increase of SLT by varying $\lambda_{\rm max}$ from the largest value ($\lambda_{\rm max}=1e10$) to the smallest one ($\lambda_{\rm max}=10$). Note that, during high traffic, the performance in terms of SLT stagnates, once $\lambda_{\rm max}$ starts crossing extremely high values. Conversely, in Fig. 4, we notice that there is a limited change of the SLT, if the value of $\lambda_{\rm max}$ varies. In fact, in low traffic scenarios, our framework aims mainly to reduce the TN energy consumption, i.e., maximising the SLT has limited importance.
To verify this analysis, Fig. 5 plots the TN energy consumption for the distinct values of $\lambda_{\rm max}$, specified previously. We remark that the TN energy consumption increases throughout the day when the value of $\lambda_{\rm max}$ decreases. Indeed, a small value of $\lambda_{\rm max}$ hampers the transmit power optimization in (23), as each update becomes negligible, making it hard to effectively drive down the energy consumption of the TN.
Through these results, we are able to grasp the trade-off between reducing energy consumption and balancing the network load efficiently. Indeed, we observe that limiting the value of $\lambda_{\rm max}$, i.e., $\lambda_{\rm max}=10$ and $\lambda_{\rm max}=1e3$, we only achieve, during high traffic hours, a minor gain in terms of network SLT of approximately $2\%$ compared to the other cases ($\lambda_{\rm max}=1e7$ and $\lambda_{\rm max}=1e10$), while the total energy consumption surges by approximately $84\%$. Considering all of this, we fix $\lambda_{\rm max}=1e7$ as this value ensures a balanced performance in both SLT and TN energy consumption.
@@
V-C Complexity and Convergence Analysis
In Fig. 6, we display the operational complexity of both proposed frameworks BLASTER and HEURISTIC as derived in Section IV-C. We notice that the operational complexity of BLASTER is higher than HEURISTIC throughout the day. This is expected, as the latter framework is implemented using less complex operations compared to the former, which resorts to multiple gradient descent methods. In fact, we see an average decrease in operational complexity of approximately $21\%$ in the day, which underlines the simplicity of HEURISTIC compared to BLASTER. Moreover, we notice that the complexity follows a pattern similar to the traffic load for both frameworks, which corroborates with the formulas derived in IV-C. Indeed, a lower number of total UEs $K$ naturally leads to a lower complexity, as seen in Fig. 6. On the opposite, a surge in traffic results in an increased complexity due to a higher number of operations needed to complete both frameworks.
To illustrate the convergence of BLASTER, we analyse the relative gain between successive iterations. A positive relative gain indicates an increase in the objective function, whereas a negative relative gain implies a decrease. Fig. 7 presents the average relative gain per iteration along with the corresponding standard deviation, computed across multiple runs of BLASTER at different hours of the day. Initially, we observe a sharp increase in relative gain, followed by a steady decline towards zero. Notably, the relative gain remains positive for the vast majority of iterations, confirming that the objective function generally improves over time. This behaviour highlights the convergence of BLASTER which on average occurs after approximately $30$ iterations.
@@ V-C Complexity and Convergence Analysis
Figure 7: Average relative gain per iteration for BLASTER.
@@ V-D Impact of the Satellite Network on Traffic Distribution
In this section, we look at the shifting role that the satellites play on the mobile network throughout the day. Fig. 8 presents the daily profile of the fraction of UEs associated with the satellite network in the different frameworks under investigation. Also, we plot the hourly number of UEs deployed in the network, represented by the black dotted line. As we have previously underlined, the priority for the proposed BLASTER and HEURISTIC frameworks during low traffic is to reduce the TN energy consumption. Hence, the satellite becomes a compelling option for offloading the TN and shutting down lightly loaded MBSs. Fig. 8 shows that BLASTER and HEURISTIC achieve a $70\,\%$ and $500\,\%$ increase of the fraction of UEs associated to the satellite network in low-traffic hours as compared to the benchmark 3GPP-NTN. Also, we notice that the proportion of UEs associated to the satellite is far greater for HEURISTIC compared to BLASTER. This is due to the fact that HEURISTIC ensures that every UE that has a signal strength greater than $RSRP_{\rm min}$ is associated to the satellite, opposed to the more sophisticated BLASTER, who would still consider the available throughput before associating to the satellite. Conversely, during high traffic hours, we notice that the satellite network takes a less prominent role for both BLASTER and HEURISTIC, essentially acting as an umbrella, providing service to UEs that would otherwise be out of coverage. Accordingly, most of of the total bandwidth is allocated to the terrestrial tier, which can efficiently serve more UEs than the non-terrestrial tier and can significantly improve the spatial reuse.